Problem 10-2 Incorrect answer. Your answer is incorrect. Try again. Selected accounts included in the property, plant, and equipment section of Stellar Corporation’s balance sheet at December 31, 2016, had the following balances. Land $426,000 Land improvements 198,800 Buildings 1,562,000 Equipment 1,363,200 During 2017, the following transactions occurred. 1. A tract of land was acquired for $213,000 as a potential future building site. 2. A plant facility consisting of land and building was acquired from Mendota Company in exchange for 28,400 shares of Stellar’s common stock. On the acquisition date, Stellar’s stock had a closing market price of $37 per share on a national stock exchange. The plant facility was carried on Mendota’s books at $156,200 for land and $454,400 for the building at the exchange date. Current appraised values for the land and building, respectively, are $326,600 and $979,800. 3. Items of machinery and equipment were purchased at a total cost of $568,000. Additional costs were incurred as follows. Freight and unloading $18,460 Sales taxes 28,400 Installation 36,920 4. Expenditures totaling $134,900 were made for new parking lots, streets, and sidewalks at the corporation’s various plant locations. These expenditures had an estimated useful life of 15 years. 5. A machine costing $113,600 on January 1, 2009, was scrapped on June 30, 2017. Double-declining-balance depreciation has been recorded on the basis of a 10-year life. 6. A machine was sold for $28,400 on July 1, 2017. Original cost of the machine was $62,480 on January 1, 2014, and it was depreciated on the straight-line basis over an estimated useful life of 7 years and a salvage value of $2,840. (a) Calculate the balance at December 31, 2017 in each of the following balance sheet accounts. (Hint: Disregard the related accumulated depreciation accounts.) Balance at December 31, 2017 Land $ Entry field with incorrect answer Land Improvements $ Entry field with incorrect answer Buildings $ Entry field with incorrect answer Equipment $ Entry field with incorrect answer
Answer #1

Working note 1 :

Common stock = 28400 * 37 = 1050800

Percentage of land = 326600 / [326600+979800] = 25%

Percentage of Building = 979800 / [326600+979800] = 75%

Allocation to land = 1050800 * 25% = 262700

Allocation to Building = 1050800 * 75% = 788100

Balance of Land at dec 31, 2017 = 426000 + 262700 = 688,700

Balance of Land improvements at dec 31, 2017 = 198800 + 134900 = 333,700

Balance of Building at dec 31, 2017 = 1562000 + 788100 = 2,350,100

Balance of Equipment at dec 31, 2017 = 1363200 + 568000 + 18460 + 28400 + 36920 -113600 - 62480 = 1,838,900